# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v3

@@ E.1 Additional Details
The mutual information for the experiments in this setting is estimated using the KL divergence between the output probability distribution of every algorithm with the prior probability, $D\left(p_{t},P^{\pi^{*}}\right)$. This is an estimator of the mutual information, similar to the one used by Seldin et al. [2011]. The mutual information was then calculated according to $D\left(p_{t},P^{\pi^{*}}\right)$, where $p_{t}$ is the real posterior. In our problem, the posterior update is done in the following manner.
@@ E.2 Additional Experiments
Figure 4: The prompting done for the LLMs in our experiments. The question is preceded by a question prompt followed by "Answer the following question:". Following this, the options are presented. <Question prompt> and <End prompt> are both replaced by a different prompt for every model.
@@ F.1 Additional details
In this experiment, we used a small LLM and a large LLM in every experiment. The small LLM was one of the following: Mistral 7B [Jiang et al., 2023], Gemma 2B [Team et al., 2024] and Falcon 7B [Almazrouei et al., 2023]. For the large LLM we used Mixtral 8x7B [Jiang et al., 2024], Llama3-70B [AI@Meta, 2024] and Gemma 7B [Team et al., 2024]. 4-bit quantization [Dettmers and Zettlemoyer, 2023] and flash attention [Dao et al., 2022] were applied for all models (excluding Falcon 7B, for which flash attention was not applied). In every round, we queried the LLM with a prompt, which is described later on. An example of the prompt used is provided in Figure 4. Using the LLM’s output we obtain a probability distribution over the tokens of the possible answers [A,B,C,D]. We then selected the answer according to this distribution and received feedback on whether the answer was correct, which was our reward. We also see that performance improvement is not clear when we use Mistral for the small model and Gemma 7B for the large one. This is because they are similar in size, while for the rest of the experiments, the larger LLM has significantly more parameters.
@@ F.3 Dataset and Models License
The MMLU dataset [Hendrycks et al., 2021a, b] is available under the MIT license. Mistral 7B [Jiang et al., 2023], Mixtral 8x7B [Jiang et al., 2024] and Falcon 7B [Almazrouei et al., 2023] are available under the Apache license 2.0. Gemma 2B and 7B [Team et al., 2024] are available under the Gemma license. Llama3-70B [AI@Meta, 2024] is available under the Llama license.

### v4

@@ E.1 Additional Details
The mutual information for the experiments in this setting is estimated using the KL divergence between the output probability distribution of every algorithm with the prior probability, $D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)$. This is an estimator of the mutual information, similar to the one used by Seldin et al. (2011). The mutual information was then calculated according to $D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)$, where $p_{t}$ is the real posterior. In our problem, the posterior update is done in the following manner.
@@ E.2 Additional Experiments
Figure 4: The prompting done for the LLMs in our experiments. The question is preceded by a question prompt followed by ”Answer the following question:”. Following this, the options are presented. <Question prompt> and <End prompt> are both replaced by a different prompt for every model.
@@ F.1 Additional details
In this experiment, we used a small LLM and a large LLM in every experiment. The small LLM was one of the following: Mistral 7B (Jiang et al., 2023), Gemma 2B (Team et al., 2024) and Falcon 7B (Almazrouei et al., 2023). For the large LLM we used Mixtral 8x7B (Jiang et al., 2024), Llama3-70B (AI@Meta, 2024) and Gemma 7B (Team et al., 2024). 4-bit quantization (Dettmers & Zettlemoyer, 2023) and flash attention (Dao et al., 2022) were applied for all models (excluding Falcon 7B, for which flash attention was not applied). In every round, we queried the LLM with a prompt, which is described later on. An example of the prompt used is provided in Figure 4. Using the LLM’s output we obtain a probability distribution over the tokens of the possible answers [A,B,C,D]. We then selected the answer according to this distribution and received feedback on whether the answer was correct, which was our reward. We also see that performance improvement is not clear when we use Mistral for the small model and Gemma 7B for the large one. This is because they are similar in size, while for the rest of the experiments, the larger LLM has significantly more parameters.
@@ F.3 Dataset and Models License
The MMLU dataset (Hendrycks et al., 2021b; a) is available under the MIT license. Mistral 7B (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024) and Falcon 7B (Almazrouei et al., 2023) are available under the Apache license 2.0. Gemma 2B and 7B (Team et al., 2024) are available under the Gemma license. Llama3-70B (AI@Meta, 2024) is available under the Llama license.